OpenStudy (nottim):

A juggler throws a ball straight up into the air with a speed of 18m/s . With what speed would she need to throw a second ball half a second later, starting from the same position as the first, in order to hit the first ball at the top of its trajectory?

OpenStudy (nottim):

So first I calculated 16.5306m, and then i calculated 1.83673s, and then gould 18.5 m/s, but apparently my answer is incorrect.

OpenStudy (nottim):

The formulas, I used, in order, were

OpenStudy (nottim):

Vf^2=Vi^2+2ad and then Vf=Vi+at and then d=V1t+1/2at^2

OpenStudy (nottim):

i got it. it was 19